OpenStudy (praetorian.10):

Find the equation of the line that passes through (0,20) and (4, 4). (You have already calculated the slope in question 13) y = -4x + 20 y = -4x+ 4 y = -0.25x + 20 y = 0.25x + 4

OpenStudy (praetorian.10):

@Nnesha

OpenStudy (zale101):

What was the slope that you calculated in question 13?

OpenStudy (praetorian.10):

-4

OpenStudy (zale101):

use the point slope formula \[y-y_1=m(x-x_1)\] plug in one of the given points in (x1,y1) and plug the slope which represents m in this formula

OpenStudy (praetorian.10):

0_o

Nnesha (nnesha):

so use this formula y=mx + b where m is a slope and b is y -intercept put the one of your favourite order pair in to the equation

OpenStudy (praetorian.10):

that looks like inter galactic alien language to me :D

OpenStudy (praetorian.10):

so y = -4x + 20

OpenStudy (praetorian.10):

no?

Nnesha (nnesha):

you know the m=-4 so you can see that c and d are not have -4 as a slope sooooo that are wrong and now we have a and b

OpenStudy (zale101):

or use y=mx+b the slope intercept form and use one of the given points, let's say (0,20) and plug them with the slope as well and substitute y=(-4)·x+b y=(-4)·(0)+b (20)=(-4)·(0)+b

Nnesha (nnesha):

right you got the right answer

OpenStudy (zale101):

you're looking for the y-intercept or b, in this equation (20)=(-4)·(0)+b

OpenStudy (praetorian.10):

thank you guys sooooooooooooo much for your patience and help :D

OpenStudy (praetorian.10):

i seriously suck

Nnesha (nnesha):

you most welcome :) :)